Project

General

Profile

NDOS Commissioning Meeting 2012/02/06

Attendance: Jon, Enrique, Martin, Mark Mathis, Athans, Leon, Jan, Susan, Nick, Minerba, Peter, Mark Messier

Agenda:

  1. Beam and Detector Update
  2. Installation Plans for the week
  3. DAQ Update and Plans for the week
  4. DCS Update and Plans for the week
  5. AOB

Beam and Detector Report

  • Beam has been down since last week, supposed to come back tomorrow.
  • In general the detector showed no significant issues this past week, although Gavin reported that he saw 0 occupancies in the nearline and event display while on shift this weekend. Peter and Jon will look into this.

Installation plans for this week

  • Mat: some testing of the APD drying system will happen this week.
  • Leon: first shipment of APDs will be small (15) and will show up ~9 days from today (silicone-coated). Two weeks later 230 will be shipped, 45 of which will be bare (which will be coated with parylene).
    • Jon: do we have all of the mounting hardware? Leon: everything except for e-boxes. Will need to make more. Need to figure out where port will be for drying system. Have enough to do 40 locations, but these will have to be modified by hand. Perhaps just the bottoms can be made now, and when the decision is made of where to put the port, then the tops can be made.
    • Jon: has the parylene company figured out how to handle the APDs without damaging them? Leon: working it out.
  • Leon: full installation will likely be completed by 2nd-3rd week of March.
  • Mark: diffusion-only drying system? Leon: yes
  • Jon: new DCMs? Peter: not clear when we'll put those on the detector.

DAQ Update and Plans for next week

  • Peter:
    • A lot of effort going on in getting FD system operational
    • More DDS testing this week (Jon: try for tomorrow morning)

DCS Update and Plans for next week

  • Athans:
    • strange error reported in DCS OPC by Gavin over the weekend. Possibly due to too many windows being open.
    • long-term planning: would like to test some hardware to be used at Ash River. Rack sensors, etc. Will need time on NDOS for this, but not until after May. Jon: no problem, since we'll have no beam at that point.

AOB

  • Jon: new ECL shift sign-up. Send constructive comments and suggestions to Jon.